OpenStudy (anonymous):

Points (3,6) and (5,-8) are endpoints of the diameter of a circle. What is the length of the diameter What is the center of the circle What is the equation of the circle

OpenStudy (phi):

do you know the distance formula ? (How to find the distance between 2 points)

OpenStudy (anonymous):

D=y2-y1 ?

OpenStudy (anonymous):

So once I calculate the distance, then what?

OpenStudy (anonymous):

2\[\sqrt{2}\] (4,-1)

OpenStudy (anonymous):

Wow you did that fast

OpenStudy (phi):

What is the center of the circle? half way between the two points. find the change in y (y2-y1), and go half way up. find the change in x (x2-x1) and go half way over.

OpenStudy (anonymous):

sdg's answer is wrong

OpenStudy (anonymous):

Thanks for actually showing how to do it phi

OpenStudy (anonymous):

I got distance is 14.14213562, the center is (2,-1)

OpenStudy (phi):

how did you get 2 in (2,-1) (the -1 looks ok)

OpenStudy (anonymous):

I used the midpoint formula. (x1+x2)/2,(y1+y2)/2

OpenStudy (phi):

well, redo the x

OpenStudy (anonymous):

Oops, 4

OpenStudy (anonymous):

(x - h)^2 + (y - k)^2 = r^2. How do I use this formula to create the equation? I know the radius, but what is h and k?

OpenStudy (phi):

The diameter is sqrt(200) (which gives you your number), but you might want to write it as sqrt(100*2)= sqrt(100)*sqrt(2)= 10sqrt(2) (that is an exact value, and sometimes on multiple choice they will use that version)

OpenStudy (phi):

(h,k) is the center of the circle.

OpenStudy (anonymous):

So I would say, (x-4)^2+(y+1)=7.07060678^2?

OpenStudy (anonymous):

(x-4)^2+(y+1)^2=7.07060678^2?

OpenStudy (phi):

ugly r! use diameter= 10sqrt(2) r= 5sqrt(2) r^2= 5*5*sqrt(2)*sqrt(2)= 25*2= 50

OpenStudy (phi):

or just square your ugly number you want r^2

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I see what you mean

OpenStudy (phi):

but that's the answer

OpenStudy (anonymous):

Thank you very much for the help!

OpenStudy (phi):

you can check if it is correct (3,6) is on the circle so (3-4)^2+(6+1)^2 should = 50 -1*-1 + 7*7= 1+49=50 yes, it works

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

Thanks a ton